### v3

# Counting spaces of functions on separable compact lines

Abstract.
We investigate the following general problem, closely related to the problem of isomorphic classification of Banach spaces $C(K)$ of continuous real-valued functions on a compact space $K$, equipped with the supremum norm:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ are there, for $K\in\mathcal{K}$?
We prove that for any uncountable regular cardinal number $\kappa$, there exist exactly $2^{\kappa}$ isomorphism types of spaces $C(K)$ for compact spaces of weight $\kappa$.
We show that, for the class $\mathcal{L}_{\omega_{1}}$ of separable compact linearly ordered spaces of weight $\omega_{1}$, the answer to the above question depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis, there are $2^{\omega_{1}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, and assuming a certain axiom proposed by Baumgartner, there is only one type.
2020 Mathematics Subject Classification:
Primary 03E35, 03E65, 03E75, 46B03, 46E15, 54F05
The second-named and the third-named authors were partially supported by the NCN (National Science Centre, Poland) research grant no. 2020/37/B/ST1/02613.
1. Introduction
One of the most fundamental and natural challenges of mathematics is the problem of classifying objects within a given class. An important part of such a challenge is an attempt to determine how many isomorphism types of objects we have in that class.
In this paper, we consider the following case of this general problem:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ of real-valued continuous functions on $K$, equipped with the supremum norm, do we have for $K\in\mathcal{K}$?
Recall that for an infinite compact space $K$, the weight of $K$ is equal to the weight and the density of $C(K)$. Hence, if the Banach spaces $C(K)$ and $C(L)$ are isomorphic, then the compact spaces $K$ and $L$ must have equal weights. It is also known that compact spaces $K$ and $L$ with isomorphic function spaces have equal cardinalities [7]. Therefore, it is natural that in the problem of counting isomorphism types of spaces $C(K)$, we restrict ourselves to the classes of compacta $K$ of fixed weight or fixed cardinality.
The classical results of Miljutin, Bessaga and Pełczyński give us a complete isomorphic classification of spaces $C(K)$ for metrizable compacta $K$, i.e., for compacta of countable weight. In particular, Bessaga and Pełczyński (cf. [34]) proved that there are exactly $\omega_{1}$ isomorphism types of $C(K)$, for countable (hence metrizable) $K$, and Miljutin (cf. [34]) showed that, for uncountable metrizable $K$, there is only one isomorphic type of $C(K)$ space. Summarizing, we have $\omega_{1}$ isomorphism types of spaces $C(K)$ for compacta $K$ of weight $\omega$ (regardless of the size of the continuum $2^{\omega}$). Recall that we have $2^{\omega}$ topological types of compacta of weight $\omega$ [24].
For the classes of nonmetrizable compacta, the chances of obtaining such complete classification results seem to be very slim, for a simple reason - we usually have too many isomorphism types of spaces $C(K)$. It is known that even if we restrict ourselves to some concrete classes of compact spaces of weight $2^{\omega}$, we still have $2^{2^{\omega}}$ isomorphism types of $C(K)$. In particular, it is true for the class of separable scattered compacta of height 3 (cf. [6]), or the class of separable, compact linearly ordered spaces (cf. [23]).
Note that, for a compact $K$ of weight $\kappa$, the cardinality of isomorphism types of $C(K)$ is bounded by $2^{\kappa}$ (cf. remarks in the last section).
In this paper, we concentrate on the case of compact spaces of weight $\omega_{1}$. We prove that there are $2^{\omega_{1}}$ isomorphism types of $C(K)$, for $K$ of such weight. Actually, we prove the following more general result:
Theorem 1.1.
Suppose that $\kappa$ is an uncountable regular cardinal. There is a family of cardinality $2^{\kappa}$ of compact spaces $K$ of weight $\kappa$, such that the corresponding Banach spaces $C(K)$ are pairwise nonisomorphic.
Most of our investigations deal with the class of separable, compact linearly ordered spaces (in short: separable compact lines). Banach spaces $C(K)$ associated with such compacta $K$ play an important role in Banach space theory. They have provided many interesting examples (see [16], [17], [18], [29]), and have been an object of investigations in numerous papers (see [2], [8], [15], [21], [23], [25], [10]).
Let $\mathcal{L}_{\kappa}$ be the class of separable, compact linearly ordered spaces of weight $\kappa$ (for our topological terminology see [11]). We show that the cardinality of isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis (i.e., $2^{\omega}=\omega_{1}$), there are $2^{2^{\omega}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$. More precisely, we prove the following:
Theorem 1.2.
If $\kappa\leq 2^{\omega}<2^{\kappa}$, then there exist $2^{\kappa}$ pairwise nonisomorphic spaces $C(L)$ for $L\in\mathcal{L}_{\kappa}$.
On the other hand, assuming a certain axiom (BA) proposed by Baumgartner, there is only one type of $C(L)$ for $L\in\mathcal{L}_{\omega_{1}}$:
Theorem 1.3 (BA).
If $K$ and $L$ are separable compact lines of weight $\omega_{1}$, then the Banach spaces $C(K)$ and $C(L)$ are isomorphic.
From this theorem we will derive the following:
Corollary 1.4 (BA).
If $K$ is a separable compact line of weight $\omega_{1}$, then the Banach space $C(K)$ is isomorphic with the direct sum $C(K)\oplus C(K)$.
The above corollary should be compared with a recent result of Kucharski [22] who proved that there exists a separable compact line $K$ of weight $2^{\omega}$ such that the Banach space $C(K)$ is not isomorphic to $C(K)\oplus C(K)$. In particular, this shows that Corollary 1.4 requires some additional set-theoretic hypothesis.
Let us recall that another example of a class $\mathcal{K}$ of compact spaces of weight $\omega_{1}$, such that the cardinality of isomorphism types of $C(K)$, for $K\in\mathcal{K}$, depends on additional set-theoretic axioms, was given in [6] ($\mathcal{K}$ is the class of separable scattered compacta of height 3).
The paper is organized as follows:
In Section 2 we explain our terminology, recall some definitions and auxiliary facts.
Section 3 contains the proof of Theorem 1.1 and its prerequisites. In Section 4 we present some general topological properties of separable compact lines, in particular counting their homeomorphic types and recalling some handy characterizations.
The main result of Section 5, dedicated to spaces of continuous functions on separable compact lines, is Lemma 5.4, which states that $C(K)$ is isomorphic to $C(K)\oplus C(M)$ for any uncountable separable compact line $K$ and any metrizable compact space $M$. Sections 6 and 7 are dedicated to proving Theorems 1.2 and 1.3, respectively.
In Section 8, we turn our attention to spaces of continuous functions on finite products of separable compact lines. Using the results of Michalak [25] and assuming the Baumgartner’s Axiom, we provide a complete isomorphic classification of such $C(K)$ spaces of weight $\omega_{1}$, see Theorem 8.5.
The last section, Section 9, is dedicated to some scattered thoughts and remarks about the subject of the paper.
2. Preliminaries
2.1. Set-theoretic terminology
We denote by $cf(\alpha)$ the cofinality of the ordinal $\alpha$. Let $\kappa$ be a regular cardinal, i.e., $cf(\kappa)=\kappa$. Recall that a set $A\subseteq\kappa$ is a club set if it is unbounded in $\kappa$ and closed in the order topology. A set $S\subseteq\kappa$ is stationary if it has a nonempty intersection with all club subsets of $\kappa$. By $\omega_{1}$ we denote the first uncountable cardinal and usually write $2^{\omega}$ for the continuum, the cardinality of the set of real numbers. In general, our set-theoretic terminology follows Jech’s book [14].
2.2. Banach spaces of continuous functions and their duals
Each compact and locally compact space we consider is Hausdorff. For a locally compact space $X$ we denote by $C_{0}(X)$ the space of real-valued continuous functions on $X$ vanishing at infinity, considered with the supremum norm. For a compact space $K$, we write $C(K)$ instead of $C_{0}(K)$, as it is the Banach space of all real-valued continuous functions on $K$. It is a classical fact that the density and weight of $C(K)$ is equal to the weight of $K$. For a set $A\subseteq K$, we denote the characteristic function of $A$ by $1_{A}$.
The dual space $C(K)^{*}$, due to the Riesz representation theorem, can be seen as $M(K)$, the Banach space of signed Radon measures on $K$ with absolute variation as the norm. There is a similar representation of $C_{0}(X)^{*}$ for any locally compact space $X$, where every element of $C_{0}(X)^{*}$ can be represented as the integration with respect to a signed Radon measure on the one-point compactification $\alpha X$ of $X$ that vanishes at the added point (see Theorem 18.4.1 of [34]).
In the space $M(K)$, equipped with the weak* topology induced by $C(K)$, there is a linearly dense topological copy of $K$, which is the set of Dirac delta measures $\Delta_{K}=\{\delta_{x}:x\in K\}$ (we will sometimes identify $K$ and $\Delta_{K}$). This, in particular, means that if the space $K$ is separable, then $M(K)$ is also separable in the weak* topology. We write $supp(\mu)$ for the support of the measure $\mu\in M(K)$.
For two Banach spaces $X,Y$ we write $X\simeq Y$ if $X$ and $Y$ are isomorphic. By $X|Y$ we mean that $Y$ contains a complemented isomorphic copy of $X$, i.e., there is a Banach space $X^{\prime}$ such that $Y\simeq X\oplus X^{\prime}$.
For a countable family of Banach spaces $\{X_{n}:n\in\omega\}$, their $c_{0}$-sum is the Banach space consisting of all sequences $(x_{n})_{n\in\omega}\in\prod_{n\in\omega}X_{n}$ such that for every $\varepsilon>0$, the set $\{n\in\omega\colon\|x_{n}\|>\varepsilon\}$ is finite, equipped with the supremum norm $\|(x_{n})_{n\in\omega}\|=\sup_{n\in\omega}\|x_{n}\|$.
The following fact is standard, see e.g. [34, Proposition 21.5.8]. Note that all hyperplanes of a Banach space are mutually isomorphic, see [12, Exercise 2.9].
Fact 2.1.
Let $K$ be a compact space with a nontrivial convergent sequence. Then the Banach space $C(K)$ is isomorphic to its hyperplanes.
In general, our terminology concerning Banach spaces follows books [12] and [34].
2.3. Scattered spaces
Recall that a topological space $F$ is scattered if every nonempty set $A\subseteq F$ contains a relative isolated point.
For a scattered space $X$, by the Cantor-Bendixson height (or simply the height) $ht(X)$ of $X$ we mean the least ordinal $\alpha$ such that the Cantor-Bendixson derivative $X^{(\alpha)}$ of the space $X$ is empty. The height of a compact scattered space is always a nonlimit ordinal.
Let us make a general remark about the bounded linear functionals on $C_{0}(X)$ for any locally compact scattered space $X$. By the Rudin theorem (Theorem 19.7.6 of [34]), if $\alpha X$ is scattered, then all Radon measures on the one-point compactification $\alpha X$ are purely atomic. From the Riesz representation theorem mentioned above it follows that any bounded linear functional $\varphi$ on $C_{0}(X)$ is defined by
$\varphi(f)=\sum_{n\in\omega}a_{n}f(x_{n}),$
where $\sum_{n\in\omega}|a_{n}|<\infty$ and $\{x_{n}:n\in\omega\}\subseteq X$ are distinct, and $f\in C_{0}(X)$.
2.4. Separable compact linearly ordered spaces
Consider an arbitrary closed subset $K$ of the unit interval $I=[0,1]$ and any subset $A\subseteq K$. Denote
$K_{A}=\big(K\times\{0\}\big)\cup\big(A\times\{1\}\big)$
and consider this set with the order topology given by the lexicographic order on $0,1]\times\{0,1\}$. If $A=(0,1)$ and $K=I$, then the space $K_{A}$ is the well-known double arrow space. Some authors use this name for the space $I_{I}$; others call the space $I_{I}$ the split interval. For any sets $A,K$ as above, the space $K_{A}$ is a separable compact line of topological weight $|A|$ (if the set $A$ is infinite), and thus it is nonmetrizable whenever the set $A$ is uncountable. If $A$ is dense in $K$, then the space $K_{A}$ is zero-dimensional. It turns out that spaces of the form $K_{A}$ exhaust the class of separable compact lines:
Theorem 2.2 (Ostaszewski, [31]).
The space $L$ is a separable compact linearly ordered space if and only if $L$ is order isomorphic (hence homeomorphic) to the space $K_{A}$ for some closed set $K\subseteq[0,1]$ and a subset $A\subseteq K$.
Due to Ostaszewski’s result, in this paper we may assume that any separable compact line is of the form $K_{A}$.
3. Counting isomorphism types of $C(K)$ of a given weight
The main result of this Section is Theorem 1.1, in which we count isomorphism types of Banach spaces of continuous functions on compact spaces of regular weight $\kappa$. Before we proceed further, let us introduce some notation. For an uncountable cardinal $\kappa$, we denote by
$E^{\kappa}_{\omega}=\{\alpha<\kappa:cf(\alpha)=\omega\}$
the set of ordinals below $\kappa$ of countable cofinality. Furthermore, if $A,B$ are any sets of ordinals, then $A<B$ means that $\alpha<\beta$ for all $\alpha\in A$, $\beta\in B$. Similarly, if $\alpha$ is an ordinal, then $\alpha<A$ ($A<\alpha$) means that $\{\alpha\}<A$ ($A<\{\alpha\}$).
The following space, known, at least for $\kappa=\omega_{1}$, as the ladder system space, seems to belong to the topological folklore since the early 1970s. Perhaps its best-known early appearances are in [9] and [33].
Definition 3.1.
Consider an uncountable regular cardinal $\kappa$. Let $\mathcal{L}=\{L_{\alpha}:\alpha\in E^{\kappa}_{\omega}\}$ be a system of ladders, i.e., for $\alpha\in E^{\kappa}_{\omega}$ we demand
  • $L_{\alpha}\subseteq\{\beta+1:\beta<\alpha\}$,
  • $L_{\alpha}\cap\beta$ is finite for all $\beta<\alpha$,
  • $\sup(L_{\alpha})=\alpha$.
For a subset $S$ of $E^{\kappa}_{\omega}$, we define the ladder system space $X_{S}(\mathcal{L})=X_{S}$ on $\kappa$:
  • The points of $X_{S}$ are elements of $\kappa$.
  • The points of $\kappa\setminus S$ are isolated in $X_{S}$.
  • The basic neighborhoods of $\alpha\in S$ are declared as all sets of the form
    $(L_{\alpha}\setminus F)\cup\{\alpha\},$
    where $F\subseteq L_{\alpha}$ is finite.
The following should be clear:
Proposition 3.2.
Suppose that $\kappa$ is an uncountable regular cardinal. Any ladder system space on $\kappa$ is a locally compact, scattered space of weight $\kappa$ and height $2$.
The one-point compactification of $X_{S}$ will be denoted $KL_{S}$. It is quite easy to prove that the space $C_{0}(X_{S})$ is isomorphic to $C(KL_{S})$.
Lemma 3.3.
Suppose that $\kappa$ is an uncountable regular cardinal and $X_{S}$ is a ladder system space on $\kappa$ for some $S\subseteq E^{\kappa}_{\omega}$. Then the Banach spaces $C_{0}(X_{S})$ and $C(KL_{S})$ are isomorphic.
Proof.
By the Stone-Weierstrass theorem, $C_{0}(X_{S})$ is a hyperplane of $C(KL_{S})$. The space $KL_{S}$ is scattered and infinite, so it admits a nontrivial convergent sequence. The rest follows from Fact 2.1. ∎
Lemma 3.4.
Suppose that $\kappa$ is an uncountable regular cardinal and $X_{R},X_{S}$ are two ladder system spaces on $\kappa$ for some sets $R,S\subseteq E^{\kappa}_{\omega}$. If $T:C_{0}(X_{R})\rightarrow C_{0}(X_{S})$ is a bounded linear operator, then for every $\gamma\in\kappa$ we have
$|\{\alpha\in X_{S}\colon T^{*}(\delta_{\alpha})(\{\gamma\})\neq 0\}|<\kappa.$
Proof.
Assume towards a contradiction that there is $\gamma\in X_{R}$ such that cardinality the set
$A=\{\alpha\in X_{S}\colon T^{*}(\delta_{\alpha})(\{\gamma\})\neq 0\}$
equals $\kappa$. By passing to a subset of $A$ of cardinality $\kappa$, we may assume that there is $\varepsilon>0$ such that $|T^{*}(\delta_{\alpha})(\{\gamma\})|>\varepsilon$ for all $\alpha\in A$. If $\gamma\in X_{R}$ is isolated in $X_{R}$, then $1_{\{\gamma\}}$ is continuous and then
$|T(1_{\{\gamma\}})(\alpha)|=|T^{*}(\delta_{\alpha})(\{\gamma\})|>\varepsilon$
for all $\alpha\in A$, which is impossible since $A$ is noncompact as a subset of $\kappa$ of cardinality $\kappa$, but continuous functions in $C_{0}(X_{S})$ must vanish at infinity.
Otherwise, $\gamma$ is not isolated in $X_{R}$. Let $L_{\gamma}$ be the ladder at $\gamma$. By passing to a subset of $A$ of cardinality $\kappa$, using the regularity of $\kappa$ and the regularity of the measure $T^{*}(\delta_{\alpha})$, we may assume that there is a cofinite subset $L$ of $L_{\gamma}$ such that $|T^{*}(\delta_{\alpha})|(L)<\varepsilon/2$ for every $\alpha\in A$. Now $1_{L\cup\{\gamma\}}$ is continuous on $X_{R}$ and
$|T(1_{L\cup\{\gamma\}})(\alpha)|=|T^{*}(\delta_{\alpha})(L\cup\gamma)|>\varepsilon-\varepsilon/2=\varepsilon/2$
for all $\alpha\in A$, which is again impossible, as continuous functions in $C_{0}(X_{S})$ must vanish at infinity. ∎
The following lemma will be the main tool for proving Theorem 1.1.
Lemma 3.5.
Suppose that $\kappa$ is an uncountable regular cardinal and that $R,S\subseteq E_{\omega}^{\kappa}$ are two sets such that $S\setminus R$ is stationary in $\kappa$. Then there is no bounded linear operator from $C_{0}(X_{R})$ into $C_{0}(X_{S})$ with a dense range.
Consequently, there is no bounded linear operator $C(KL_{R})$ into $C(KL_{S})$ with a dense range.
Proof.
Suppose that $T:C_{0}(X_{R})\rightarrow C_{0}(X_{S})$ is a bounded linear operator. We will show that it does not have a dense range. Consider the adjoint operator $T^{*}:C_{0}(X_{S})^{*}\rightarrow C_{0}(X_{R})^{*}$. Let $F,G:\kappa\rightarrow\kappa$ be nondecreasing functions such that
$supp(T^{*}(\delta_{\alpha}))\subseteq F(\alpha),$
$\bigcup\{supp(T^{*}(\delta_{\beta})):G(\alpha)<\beta<\kappa\}\cap\alpha=\emptyset$
for every $\alpha<\kappa$. The existence of $F$ follows from the fact that the supports of Radon measures on $\alpha X_{R}$ that vanish on the one point added in the compactification are countable and such sets must be bounded in $\kappa$ since it is an uncountable regular cardinal.
The proof of the existence of $G$ uses the fact that $T$ is bounded: if the value $G(\alpha)$ could not be found, there would exist $A\subseteq\kappa$ of cardinality $\kappa$ and a fixed $\gamma<\alpha$ such that $T^{*}(\delta_{\beta})(\{\gamma\})\not=0$ for all $\beta\in A$, which contradicts Lemma 3.4.
Using the standard closure argument one can easily verify that, for any function $H:\kappa\rightarrow\kappa$, the set $\{\alpha<\kappa\colon H[\alpha]\subseteq\alpha\}$ is a club set in $\kappa$. Since the intersection of two club sets in $\kappa$ is a club set in $\kappa$ we obtain a club set $C\subseteq\kappa$ which is closed under $F$ and $G$, that is, for every $\alpha\in C$ we have $F[\alpha],G[\alpha]\subseteq\alpha$.
Fix $\alpha\in C\cap(S\setminus R)$. By recursion, construct a strictly increasing sequence $\beta_{n}^{\prime}\in L_{\alpha}$ such that $F(\beta_{n}^{\prime}),G(\beta_{n}^{\prime})<\beta_{n+1}^{\prime}$. This can be done since $C$ is closed under $F$ and $G$, and $L_{\alpha}$ is cofinal in $\alpha$. The definitions of $F$ and $G$ imply that
$supp(T^{*}(\delta_{\beta_{n+1}^{\prime}}))\subseteq[\beta_{n}^{\prime},\beta_{n+2}^{\prime})$
for every $n\in\omega$.
Now define $\beta_{n}=\beta_{2n}^{\prime}$ and note that the supports of $T^{*}(\delta_{\beta_{n}})$ satisfy
$supp(T^{*}(\delta_{\beta_{n}}))<\beta_{2n+1}^{\prime}<supp(T^{*}(\delta_{\beta_{n+1}}))<\alpha.$
We can see that $(\beta_{n})_{n\in\omega}$ converges to $\alpha$ in $X_{S}$ as it is a strictly increasing sequence included in $L_{\alpha}$. Hence $(\delta_{\beta_{n}})_{n\in\omega}$ weakly converges to $\delta_{\alpha}$ in $C_{0}(X_{S})^{*}$. The weak continuity of $T^{*}$ yields the weak convergence of $(T^{*}(\delta_{\beta_{n}}))_{n\in\omega}$ to $T^{*}(\delta_{\alpha})$ in $C_{0}(X_{R})^{*}$.
We will show that in fact $(T^{*}(\delta_{\beta_{n}}))_{n\in\omega}$ weakly converges to 0 in $C_{0}(X_{R})$, which will imply $T^{*}(\delta_{\alpha})=0$ showing that $T^{*}$ is not injective and hence $T$ cannot have a dense range.
Indeed, when $f\in C_{0}(X_{R})$, then for every $\varepsilon>0$ there is $\beta<\alpha$ such that all values of $f|[\beta,\alpha)$ are in the interval $(-\varepsilon,\varepsilon)$. Otherwise, there would be a strictly increasing and cofinal sequence $(\alpha_{n})_{n\in\omega}$ of points in $\alpha$ satisfying $|f(\alpha_{n})|\geq\varepsilon$. As the point $\alpha$ is isolated in $X_{R}$ and the basic neighborhood of other points in $X_{R}$ have bounded intersections with $\alpha$, the sequence $(\alpha_{n})_{n\in\omega}$ is divergent to infinity, which is a contradiction.
It follows from the above and the fact that $T^{*}(\delta_{\beta_{n}})$ form a bounded sequence of measures that the sequence $T^{*}(\delta_{\beta_{n}})(f)$ converges to 0 for each $f\in C_{0}(X_{R})$. Thus, we have that the sequence $(T^{*}(\delta_{\beta_{n}}))_{n\in\omega}$ weakly converges to 0 in $C_{0}(X_{S})^{*}$ and $T^{*}(\delta_{\alpha})=0$ as needed to conclude that $T$ does not have dense range. ∎
We can now proceed to the proof of Theorem 1.1.
Proof of Theorem 1.1.
Fix $\kappa$ as in the statement. By a theorem of Solovay (Theorem 8.10 of [14] also cf. Lemma 8.8 of [14]) there is a pairwise disjoint family $\{S_{\xi}:\xi<\kappa\}$ of stationary subsets of $E_{\omega}^{\kappa}$. For every $A\subseteq\kappa$ consider
$S(A)=\bigcup_{\xi\in A}S_{\xi}.$
Note that $S(A)\triangle S(B)$ is stationary for any two distinct $A,B\subseteq\kappa$. So, by Lemmas 3.5 and 3.3 the family $\{C(KL_{S(A)})\colon A\subseteq\kappa\}$ is the desired family of Banach spaces. ∎
4. Topological properties of separable compact lines
In this section we use our convention that we identify separable compact lines with the spaces of the form $K_{A}$, where $K\subseteq[0,1]$ and $A\subseteq K$, see subsection 2.4.
First, we will show that for any cardinal number $\kappa$ satisfying $\omega\leq\kappa\leq 2^{\omega}$, there are $2^{\kappa}$ pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$ (recall that, for infinite $A$, the weight of $K_{A}$ is equal to $|A|$).
We start with the following observation of van Douwen [37, Section 10].
Proposition 4.1.
Let $K_{A},L_{B}$ be separable compact lines that are homeomorphic. Then there exist countable sets $C\subseteq K$ and $D\subseteq L$ such that $K\setminus(A\cup C)$ and $L\setminus(B\cup D)$ are homeomorphic.
The above proposition together with a standard fact that each subset of $I$ can be homeomorphic to at most continuum many subsets of $I$, and some routine cardinal calculations, easily implies that for a fixed separable compact line $K_{A}$ we have at most $2^{\omega}$ many separable compact lines $L_{B}$ homeomorphic to $K_{A}$ (cf. [37, Section 10], [23, Section 2]). This immediately allows us to obtain the following variation of van Douwen’s result that there are $2^{2^{\omega}}$ separable compact lines.
Corollary 4.2.
If a cardinal number $\kappa\leq 2^{\omega}$ satisfies $2^{\kappa}>2^{\omega}$, then there are $2^{\kappa}$ pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$.
We will show that the above result also holds true when $2^{\kappa}=2^{\omega}$.
Theorem 4.3.
For any infinite cardinal number $\kappa\leq 2^{\omega}$, there are $2^{\kappa}$ pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$.
Proof.
By Corollary 4.2 it is enough to consider the case when $2^{\kappa}=2^{\omega}$, and construct a family $\mathcal{K}$ of size $2^{\omega}$ consisting of pairwise nonhomeomorphic spaces $K_{A}$ of weight $\kappa$.
By a classical result of Mazurkiewicz and Sierpiński [24] there is a family $\mathcal{L}$ of size $2^{\omega}$ consisting of pairwise nonhomeomorphic compact subsets of the interval $[0,1/3]$. If $\kappa=\omega$, then we are done. Otherwise, we take a subset $A$ of $(2/3,1)$ such that $|A\cap P|=\kappa$, for any nonempty open subinterval $P$ of $(2/3,1)$, and we put $\mathcal{K}=\{L\cup[2/3,1]_{A}:L\in\mathcal{L}\}$. The family $\mathcal{K}$ has the required properties, since each point of $2/3,1]_{A}$ has only nonmetrizable neighborhoods, hence $L\cup[2/3,1]_{A}$ and $L^{\prime}\cup[2/3,1]_{A}\in\mathcal{K}$ are homeomorphic if and only if the spaces $L$ and $L^{\prime}$ are homeomorphic. ∎
Proposition 4.4.
Each separable compact line is hereditarily separable, hereditarily Lindelöf, and Fréchet.
Proof.
Every separable compact line has to be Fréchet, since it is a first countable space. Each separable compact line is of the form $K_{A}$, so it is a continuous image of the double arrow space $I_{(0,1)}$ and all these properties are preserved by continuous images. The double arrow space is hereditarily separable, hereditarily Lindelöf, as it is a union of two disjoint copies of the Sorgenfrey line, which is such. ∎
Observe that the above proposition implies, in particular, that the class of separable compact lines is closed under taking closed subsets.
[…]
Corollary 4.5.
Let $L$ be a separable compact line and $\alpha=\min\{\beta:L^{(\beta+1)}=L^{(\beta)}\}$. Then $\alpha<\omega_{1}$ and $|L\setminus L^{(\alpha)}|\leq\omega$.
If a separable compact line is perfect (without isolated points), we can modify it by a homeomorphism to a more convenient form.
Proposition 4.6.
Let $L$ be a separable compact line. If $L$ has no isolated points, then there exists $B\subseteq(0,1)$ such that $L$ is order isomorphic (hence homeomorphic) to $I_{B}$.
Proof.
Without loss of generality, we can assume that $L=K_{A}$, where $A\subseteq K\subseteq I$, and $K$ is a closed set such that $\inf K=0$ and $\sup K=1$.
Enumerate the family of all connectedness components of $I\setminus K$ as $\{(a_{n},b_{n}):n\in N\}$, where $N=\omega$ or $N\in\omega$. Observe that, for each $n\in N$, the set $\{a_{n},b_{n}\}$ is disjoint from $A$, since $K_{A}$ has no isolated points.
Now consider an equivalence relation on $K$ defined by $a_{n}\!\sim b_{n}$ for all $n\in N$ and let $q:K\to K/\!\!\sim$ be the quotient map. Notice that the relation $\sim$ is a closed relation (see [11]), as every increasing or decreasing sequence of $a_{n}$’s will have the same limit as a sequence of $b_{n}$’s with the same indicies. One can easily check that the resulting quotient space $K/\!\!\sim$ is a linearly ordered connected metrizable compact space, hence it is order isomorphic to the unit interval (cf. [11, 6.3.2]). Denote by $h:K/\!\!\sim\ \to I$ the order isomorphism satisfying $h(q(0))=0$. One can verify that $K_{A}$ can be identified with $I_{B}$ where $B=h[q[A\cup\{a_{n}:n\in N\}]]$. ∎
Proposition 4.7.
Let $L$ be a separable zero-dimensional compact line. Let $C\subseteq I$ be the ternary Cantor set. If $L$ has no isolated points, then there exists $B\subset C$ such that $L$ is order isomorphic (hence homeomorphic) to $C_{B}$.
Proof.
By Proposition 4.6, we can assume that $L=I_{A}$ for some $A\subseteq I$. As $L$ is zero-dimensional, $A$ is dense in $I$. We can find a copy of the dyadic numbers $Q=\{a_{s}:s\in 2^{<\omega}\}$ dense in $A$, so that $a_{s}<a_{s^{\prime}}$ (in $A$) if and only if $s<s^{\prime}$ (in $2^{<\omega}$ with the lexicographical order). Let $\{(q_{s}^{0},q_{s}^{1})\subseteq I:s\in 2^{<\omega}\}$ be the usual, order preserving enumeration of the intervals removed during the construction of the Cantor set $C\subseteq I$.
Define a function $h:Q\times\{0,1\}\to C$ by
$h(a_{s},b)=q_{s}^{b}$
for $s\in 2^{<\omega},b\in\{0,1\}$. Since $h$ is order-preserving (with respect to the lexicographic order on $Q\times\{0,1\}$), $Q$ is dense in $I$, and $\{q_{s}^{0},q_{s}^{1}:s\in 2^{<\omega}\}$ is dense in $C$, we can extend $h$ to an order isomorphism $\widetilde{h}:I_{A}\to C_{B}$, where $B\times\{1\}=\widetilde{h}\big[(A\setminus Q\big)\times\{1\}]$. ∎
5. General properties of spaces $C(K_{A})$
For a closed subset $F$ of a compact space $K$, we denote by $C_{0}(K\|F)$ the subspace $\{f\in C(K):f|F\equiv 0\}$ of $C(K)$.
Proposition 5.1.
Let $F$ be a closed subset of a Fréchet compact space $K$, such that $K\setminus F$ is infinite. Then $C_{0}(K\|F)\simeq C(K/F)$.
Proof.
The space $C_{0}(K\|F)$ is isomorphic to a hyperplane of $C(K/F)$. From our hypothesis, it follows that $K/F$ contains a nontrivial convergent sequence, thus the statement follows from Fact 2.1. ∎
Corollary 5.2.
Let $F$ be a closed subset of a separable compact line $K$. Then there is a compact space $L$ such that $C_{0}(K\|F)\simeq C(L)$.
Proof.
If $K\setminus F$ is infinite, then we can apply the above proposition. Otherwise, the space $C_{0}(K\|F)$ is finite-dimensional, and we can take as $L$ a discrete finite space of appropriate size. ∎
The next result follows from a theorem of Heath and Lutzer [13, Theorem 2.4] which provides the existence of an extension operator $T:C(L)\to C(K)$ for compact lines $L\subseteq K$ (see [23, Lemma 4.3] for a self-contained proof).
Lemma 5.3 (Marciszewski, [23]*Lemma 4.3).
Let $K_{A}$ be a separable compact line. For each nonempty closed subset $F$ of $K_{A}$, the space $C(K_{A})$ is isomorphic to $C(F)\oplus C_{0}(K_{A}\|F)$.
The following result is a more general version of [23, Lemma 4.6]. Note that it is true in ZFC, without any additional set-theoretic hypothesis. The proof of this lemma follows closely the argument in [23], we present it here for the convenience of readers.
Lemma 5.4.
For every uncountable separable compact line $K_{A}$ and every nonempty metrizable compact space $M$ the spaces $C(K_{A})$ and $C(K_{A})\oplus C(M)$ are isomorphic.
Proof.
First, let us show that $C(2^{\omega})|C(K_{A})$.
As $K$ is an uncountable closed subset of the unit interval, we can find a copy of the Cantor set $C\subseteq K$. Put $B=C\cap A$ and see that $C_{B}$ is a closed subspace of $K_{A}$. By Lemma 5.3, we have $C(C_{B})|C(K_{A})$. Now, it is sufficient to show that $C(2^{\omega})|C(C_{B})$.
Write $p_{1}:C_{B}\to C$ for the projection. Fix a homeomorphism $h:C\to C_{1}\times C_{2}$, where both $C_{1},C_{2}$ are copies of $2^{\omega}$ and by $\pi_{1},\pi_{2}$ denote the projections from $C_{1}\times C_{2}$ onto $C_{1},C_{2}$, respectively.
Our goal is to construct a regular averaging operator for the map
$\varphi=\pi_{1}\circ h\circ p_{1}:C_{B}\to C_{1},$
i.e., a positive linear operator $T:C(C_{B})\to C(C_{1})$ satisfying $T(1_{C_{B}})=1_{C_{1}}$ and $T(g\circ\varphi)=g$ for every $g\in C(C_{1})$. The existence of such an operator implies that $C(C_{1})|C(L)$, see [32, Proposition 8.2].
The set $C_{2}$ is a copy of $2^{\omega}$, so there is a standard product measure on $C_{2}$, denote it by $\mu$. Define the operator $T:C(C_{B})\to C(C_{1})$ by the formula
$T(f)(x)=\int_{C_{2}}f(h^{-1}(x,y),0)\mathop{}\!\mathrm{d}\mu(y)$
for $f\in C(C_{B})$ and $x\in C_{1}$. It is a well-known fact that, for every function $f\in C(C_{B})$, the function $f_{0}(t)=f(t,0)$ has at most countably many points of discontinuity (see e.g. [23, Proof of Thm. 3.7]) and is therefore measurable. It follows that $T(f)(x)$ is well defined for every $x\in C_{1}$, it is also clear that $T$ is positive and $T(1_{C_{B}})=1_{C_{1}}$. Additionally, for every $g\in C(C_{1})$, $x\in C_{1}$ and $y\in C_{2}$ we have $g\circ\varphi(h^{-1}(x,y),0)=g(x)$, hence $T(g\circ\varphi)=g$.
It remains to check that $T(f)$ is continuous on $C_{1}$ for every $f\in C(C_{B})$. Consider any convergent sequence $x_{n}\to x$ in $C_{1}$. Then for every $y\in C_{2}$ we have $h^{-1}(x_{n},y)\to h^{-1}(x,y)$, as $h$ is a homeomorphism. Consider again the function $f_{0}(t)=f(t,0)$ and let $P$ be the set of points of discontinuity of $f_{0}$ ($P$ is countable). Denote $f_{n}(y)=f_{0}(h^{-1}(x_{n},y))$ and $\widetilde{f}(y)=f_{0}(h^{-1}(x,y))$. If $R=\pi_{2}\circ h(P)$, then for every point $y\in C_{2}\setminus R$ we have $f_{n}(y)\to\widetilde{f}(y)$, thus the sequence $(f_{n})_{n\in\omega}$ is $\mu$-a.e. convergent to $\widetilde{f}$ on $C_{2}$ (note that $\mu(R)=0$, as $R$ is countable and the measure $\mu$ is nonatomic). As all functions $|f_{n}|$ are bounded by the norm of $f$, from the Lebesgue dominated convergence theorem it follows that $T(f)(x_{n})\to T(f)(x)$.
Now consider any nonempty metrizable compact space $M$. It is a well-known fact that $C(M)\oplus C(2^{\omega})\simeq C(2^{\omega})$ (e.g. by Miljutin’s theorem). Since $C(2^{\omega})$ is a factor of $C(K_{A})$, it follows that $C(M)\oplus C(K_{A})\simeq C(K_{A})$, as
$C(M)\oplus C(K_{A})\simeq C(M)\oplus C(2^{\omega})\oplus E\simeq C(2^{\omega})\oplus E\simeq C(K_{A}),$
for $E$ such that $C(K_{A})\simeq C(2^{\omega})\oplus E$. ∎
Corollary 5.5.
Let $L$ be an uncountable separable compact line and $\alpha=\min\{\beta:L^{(\beta+1)}=L^{(\beta)}\}$. Then the spaces $C(L)$ and $C(L^{(\alpha)})$ are isomorphic.
Proof.
By Corollary 4.5 we have $|L\setminus L^{(\alpha)}|\leq\omega$. Let $K$ be a compact space given by Corollary 5.2 such that $C_{0}(L\|L^{(\alpha)})\simeq C(K)$. Then either $K=L/L^{(\alpha)}$ or $K$ is finite, in both cases the space $K$ is metrizable since $L\setminus L^{(\alpha)}$ is countable. Therefore, we can use Lemmas 5.3 and 5.4 to conclude that
$C(L)\simeq C(L^{(\alpha)})\oplus C_{0}(L\|L^{(\alpha)})\simeq C(L^{(\alpha)})\oplus C(K)\simeq C(L^{(\alpha)}).\qed$
The next theorem follows immediately from the corollary above and Proposition 4.6.
Theorem 5.6.
Let $L$ be an uncountable separable compact line. Then there exists $A\subseteq(0,1)$ such that $C(L)\simeq C(I_{A})$.
6. Counting spaces $C(K_{A})$ of weight $\kappa$ when $2^{\kappa}>2^{\omega}$
It is a well-known fact that, consistently (e.g. under CH), $2^{\omega_{1}}$ can be greater than $2^{\omega}$. The purpose of this Section is the proof of Theorem 1.2, which we will derive from a result obtained in [6].
First, we count the number of potential isomorphisms between spaces $C(K_{A})$. Let us recall the following, very general theorem due to Cabello-Sánchez, Castillo, Marciszewski, Plebanek and Salguero-Alarcón.
Theorem 6.1.
[6, Theorem 5.1] Let $\mathcal{K}$ be a family of compact spaces such that
  • $K$ is separable and $|M(K)|=2^{\omega}$ for every $K\in\mathcal{K}$;
  • For every pair of distinct $K,L\in\mathcal{K}$ one has $C(K)\simeq C(L)$ and $K,L$ are not homeomorphic.
Then $\mathcal{K}$ is of cardinality at most $2^{\omega}$.
Proof of Theorem 1.2.
It is a known fact that for any separable compact line $K_{A}$ we have $|M(K_{A})|=2^{\omega}$, see [18].
Due to Theorem 4.3, there exists a family $\mathcal{K}$ of size $2^{\kappa}$ consisting of pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$. For a fixed $K_{A}\in\mathcal{K}$, by Theorem 6.1, we have at most $2^{\omega}$ spaces $L_{B}\in\mathcal{K}$ such that $C(K_{A})\simeq C(L_{B})$. Now, the statement of Theorem 1.2 easily follows from the hypothesis that $2^{\kappa}>2^{\omega}$. ∎
7. Counting spaces $C(K_{A})$ under Baumgartner’s Axiom
In this Section we will prove Theorem 1.3, i.e., we will show that consistently we can have a situation completely opposite to the one described by Theorem 1.2. The proof is presented at the end of this section, after all the necessary lemmas.
Note that by Theorem 4.3 there exist (in ZFC) $2^{\omega_{1}}$ pairwise nonhomeomorphic separable compact lines of weight $\omega_{1}$.
Classical Cantor’s isomorphism theorem states that all countable dense subsets of $\mathbb{R}$ are order isomorphic. The same cannot be said about uncountable subsets, but, after adding a very natural condition, a similar statement can be proved to be consistent with ZFC. For formulating this statement we need the following definition.
Definition 7.1.
Let $\kappa$ be an infinite cardinal. We say that a subset $A$ of a topological space $X$ is $\kappa$-dense in $X$, if for any nonempty open subset $U$ of $X$ we have $|U\cap A|=\kappa$.
We will mainly use this definition in the case when $X$ is a subspace of the real line $\mathbb{R}$, so we will consider only cardinals $\kappa\leq 2^{\omega}$. Obviously, a subset $A$ of $\mathbb{R}$ is $\kappa$-dense in $\mathbb{R}$ if, for every nonempty open interval $I\subseteq\mathbb{R}$, we have $|I\cap A|=\kappa$. We refer to the following statement about $\kappa$-dense sets in $\mathbb{R}$ as Baumgartner’s Axiom.
BA($\kappa$): Every two sets of reals that are $\kappa$-dense in $\mathbb{R}$ are order isomorphic.
This axiom was introduced as a natural consequence of a classical theorem of Baumgartner [3], who proved that consistently all $\omega_{1}$-dense sets of reals are order isomorphic, i.e., using the above language he showed the following (later on, we will denote $\textsf{BA}(\omega_{1})$ simply as BA).
Theorem 7.2 (Baumgartner).
[…]
On the other hand, it should be clear that $\textsf{BA}(\mathfrak{c})$ is false, so consistently $\textsf{BA}(\omega_{1})$ is false (e.g. under CH). Todorčević in [36] even showed that $\textsf{BA}(\mathfrak{b})$ is also false. Theorem 7.2 can be proven by forcing (as in the original proof), or using PFA as in [4, Theorem 6.9, p. 945].
The consistency of $\textsf{BA}(\omega_{2})$ is an important open problem in set theory, see [28].
Using Dedekind cuts, one can easily obtain the following lemma.
Lemma 7.3.
Every order isomorphism $f:A\to B$ between $\omega_{1}$-dense sets $A,B\subseteq\mathbb{R}$ extends to an order isomorphism $\widetilde{f}$$\mathbb{R}\to\mathbb{R}$.
We can easily transfer such results from the real line to the unit interval $(0,1)$, since these two spaces are order isomorphic. Therefore, we can use Lemma 7.3 to prove some structural results about separable compact lines.
Corollary 7.4.
If $\omega_{1}$-dense sets $A,B\subseteq(0,1)$ are order isomorphic, then the spaces $I_{A},I_{B}$ are homeomorphic.
Proof.
Suppose that there exists an order isomorphism $f:A\to B$. By Lemma 7.3 and the remark following it, we extend it to an order isomorphism $\widetilde{f}:(0,1)\to(0,1)$, which in turn, can be extended in an obvious way to an order isomorphism $\hat{f}:[0,1]\to[0,1]$. Consider the map $\varphi:I_{A}\to I_{B}$ defined by the formula $\varphi(x,i)=(\hat{f}(x),i)$ for $(x,i)\in I_{A}$. Clearly, $\varphi$ is an order isomorphism between $I_{A}$ and $I_{B}$, equipped with the lexicographic orders, hence it is a homeomorphism. ∎
As a consequence, it should be clear that under BA we obtain the following.
Corollary 7.5 (BA).
If sets $A,B\subseteq(0,1)$ are $\omega_{1}$-dense, then the spaces $I_{A}$ and $I_{B}$ are homeomorphic.
We need one more lemma which does not require any additional set-theoretic axioms.
Lemma 7.6.
For any separable compact line $K$ of weight $\omega_{1}$, the space $C(K)$ is isomorphic to a space $C(I_{B})$ for some $\omega_{1}$-dense set $B\subseteq(0,1)$.
Proof.
Using Theorem 5.6 we can assume that $K=I_{A}$, where $A$ is a subset of $(0,1)$ of size $\omega_{1}$. Let
$U=\bigcup\{(a,b)\subseteq(0,1):a,b\in\mathbb{Q}\mbox{ and }|(a,b)\cap A|\leq\omega\}.$
Clearly, $|U\cap A|\leq\omega$ and $A\setminus U$ is $\omega_{1}$-dense in a compact set $M=I\setminus U$. We will consider a separable compact line $L=M_{A\setminus U}$. First, note that $L$ is a closed subset of $K$, hence, by Lemma 5.3 we have
(1) $C(K)\simeq C(L)\oplus C_{0}(K\|L)\,.$
Next, let us verify that the space $C_{0}(K\|L)=C_{0}(I_{A}\|M_{A\setminus U})$ is separable.
The space $C(I_{A\cap U})$ is separable, since $A\cap U$ is countable; therefore its subspace $C_{0}(I_{A\cap U}\|M\times\{0\})$ is also separable. Let $\varphi:I_{A}\to I_{A\cap U}$ be a natural continuous surjection defined by the formula:
$\varphi((x,i))=\begin{cases}(x,0)&\quad\mbox{for }(x,i)\in(A\setminus U)\times\{1\}\,,\\ (x,i)&\quad\mbox{for }(x,i)\in I_{A}\setminus[(A\setminus U)\times\{1\}]\,.\end{cases}$
One can easily check that the isometric embedding $T_{\varphi}:C(I_{A\cap U})\to C(I_{A})$ generated by $\varphi$, i.e., an operator sending each $f\in C(I_{A\cap U})$ to the composition $f\circ\varphi$, maps the space $C_{0}(I_{A\cap U}\|M\times\{0\})$ onto $C_{0}(I_{A}\|M_{A\setminus U})$, hence the latter space is also separable.
By Corollary 5.2 the space $C_{0}(K\|L)=C_{0}(I_{A}\|M_{A\setminus U})$ is isomorphic to the space $C(N)$ for some compact space $N$. Since $C_{0}(K\|L)$ is separable, $N$ must be metrizable, so we can use Lemma 5.4 and property (1) to conclude that $C(K)$ and $C(L)$ are isomorphic.
Let $\alpha=\min\{\beta:L^{(\beta+1)}=L^{(\beta)}\}$. By Corollary 4.5 we have $|L\setminus L^{(\alpha)}|\leq\omega$. The space $L^{(\alpha)}$ has no isolated points, therefore, by Proposition 4.6 it is order isomorphic (hence homeomorphic) to a space $I_{B}$ for some $B\subseteq(0,1)$.
The set $A\setminus U$ is $\omega_{1}$-dense in $M$, hence each nonempty open interval in $L=M_{A\setminus U}$ (an interval with respect to lexicographic order in $L$) contains $\omega_{1}$ many gaps, i.e., empty intervals of the form $\big((a,0),(a,1)\big)$, where $a\in A\setminus U$. Since $|L\setminus L^{(\alpha)}|\leq\omega$, the spaces $L^{(\alpha)}$ and $I_{B}$ have the same property, that is, every nonempty order open interval contains $\omega_{1}$ gaps. This means that the set $B$ is $\omega_{1}$-dense in $I$. By Corollary 5.5 the spaces $C(L)$ and $C(L^{(\alpha)})$ are isomorphic, hence we obtain the desired conclusion
$C(K)\simeq C(L)\simeq C(L^{(\alpha)})\simeq C(I_{B})\,.\qed$
We can now complete the proof of the main result of this section.
Proof of Theorem 1.3.
The theorem follows immediately from Corollary 7.5 and Lemma 7.6. ∎
Corollary 1.4 easily follows from Theorem 1.3:
Proof of Corollary 1.4.
Assume BA, and let $K$ be an arbitrary separable compact line of weight $\omega_{1}$. The direct sum $C(K)\oplus C(K)$ can be identified with the space $C(K\sqcup K)$, where $K\sqcup K$ is a disjoint union of two copies of the space $K$. Since $K\sqcup K$ is homeomorphic to a separable compact line of weight $\omega_{1}$, Theorem 1.3 implies that $C(K)$ and $C(K)\oplus C(K)$ are isomorphic. ∎
8. Function spaces on finite products of separable compact lines
In this section, using Theorem 1.3, the results of Michalak from [25], and assuming Baumgartner’s Axiom, we consistently obtain a complete isomorphic classification of spaces $C(K)$, where $K$ is a nonmetrizable finite product of separable compact lines of weight $\leq\omega_{1}$.
For the entirety of this Section, fix some set $\mathbf{A}\subseteq(0,1)$ of cardinality $\omega_{1}$. Define a family of compact spaces
$\mathcal{K}_{\mathbf{A}}=\{(I_{\mathbf{A}})^{n}:n\geq 1\}$ $\cup\{(I_{\mathbf{A}})^{n}\times[0,1]:n\geq 1\}$
$\cup\{(I_{\mathbf{A}})^{n}\times(\omega^{\omega^{\alpha}}+1):n\geq 1,\omega_{1}>\alpha>0\}.$
The goal of this Section is to prove that the class $\mathcal{K}_{\mathbf{A}}$ (regardless of which set $\mathbf{A}\subseteq(0,1)$ of cardinality $\omega_{1}$ is chosen) exhausts all isomorphism types of $C(K)$, where $K$ is any nonmetrizable finite product of separable compact lines of weight $\leq\omega_{1}$. First, let us note that the spaces $C(K)$ for $K\in\mathcal{K}_{\mathbf{A}}$ are pairwise nonisomorphic due to the following result of Michalak.
Theorem 8.1 ([25]*Corollary 5.3).
Let $n,k\in\mathbb{N}$. Let $K_{1},\dotsc,K_{n},L_{1},\dotsc,L_{k}$ be nonmetrizable separable compact lines.
  1. (a)
    If $M_{1},M_{2}$ are infinite compact metric spaces such that the spaces $C(\prod_{i=1}^{n}K_{i}\times M_{1})$ and $C(\prod_{i=1}^{k}L_{i}\times M_{2})$ are isomorphic, then $n=k$ and the spaces $C(M_{1})$ and $C(M_{2})$ are isomorphic.
  2. (b)
    If $M$ is a compact metric space such that the spaces $C(\prod_{i=1}^{n}K_{i}\times M)$ and $C(\prod_{i=1}^{k}L_{i})$ are isomorphic, then $n=k$ and the space $C(M)$ is finite-dimensional or isomorphic to $c_{0}$.
Let us recall some notation and simple facts regarding the spaces of vector-valued functions. For a compact space $K$ and a Banach space $(X,\|\cdot\|_{X})$, by $C(K,X)$ we denote the space of continuous functions $f:K\to X$ with the supremum norm $\|f\|=\sup_{k\in K}\|f(k)\|_{X}$.
Fact 8.2 ([34]*Proposition 7.7.5).
If $K,L$ are compact spaces, then $C(K\times L)\simeq C(K,C(L))$.
The next fact is obvious.
Fact 8.3.
For any compact space $K$ and isomorphic Banach spaces $X,Y$, we have $C(K,X)\simeq C(K,Y)$.
To fully classify the spaces of continuous functions on compact spaces of the family $\mathcal{K}_{\mathbf{A}}$, we also need the following result.
Lemma 8.4 (BA).
If a set $A$ is $\omega_{1}$-dense in $(0,1)$, then the Banach spaces $C(I_{A})$ and $C(I_{A},c_{0})$ are isomorphic.
Proof.
First, recall a simple observation that, for any compact space $K$, the space $C(K,c_{0})$ is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$. Indeed, $c_{0}$ is isometric to $C_{0}(\omega)$, hence $C(K,c_{0})$ is isometric to $C(K,C_{0}(\omega))$, which in turn is isometric to $C_{0}(K\times\omega)$, see [34]*Proposition 7.7.6. Clearly, this latter space is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$.
Let $a_{0}=1$ and $(a_{n})_{n\geq 1}$ be a decreasing sequence in $A$ which converges to 0. Put
$K_{n}=[(a_{n+1},1),(a_{n},0)]\subseteq I_{A}$
for $n\in\omega$ (an interval with respect to lexicographic order in $I_{A}$). For every $n\in\omega$, the set $A\cap K_{n}$ is $\omega_{1}$-dense in $(a_{n+1},a_{n})$, so by Corollary 7.5, $C(K_{n})$ is isometric to $C(I_{A})$. As $I_{A}$ is Fréchet, due to Fact 2.1, we have $C(I_{A})\simeq C_{0}(I_{A}\|\{(0,0)\})$.
What remains is a standard observation that $C_{0}(I_{A}\|\{(0,0)\})$ is isometric to the $c_{0}$-sum of spaces $C(K_{n}),n\in\omega$. ∎
Now we have all the tools to prove the main result of this Section.
Theorem 8.5 (BA).
Let $K$ be a nonmetrizable finite product of separable compact lines of weight $\leq\omega_{1}$. Then the space $C(K)$ is isomorphic to exactly one space $C(L)$ for $L\in\mathcal{K}_{\mathbf{A}}$.
Proof.
By Theorem 8.1, for two distinct spaces $K_{1},K_{2}\in\mathcal{K}_{\mathbf{A}}$, we have $C(K_{1})\not\simeq C(K_{2})$. Thus, it is enough to prove that, for any $m\geq 1$, and any separable compact lines $K_{1},\dotsc,K_{m}$ of weight $\leq\omega_{1}$, if at least one of them is nonmetrizable, then there is some $L\in\mathcal{K}_{\mathbf{A}}$ such that $C(\prod_{i=1}^{m}K_{i})\simeq C(L)$.
Without loss of generality, assume that for some $1\leq n\leq m$, we have $w(K_{i})=\omega_{1}\iff i\leq n$. Then $M=\prod_{i=n+1}^{m}K_{i}$ is a metrizable compact space.
By Facts 8.2, 8.3 we have
$C\left(\prod_{i=1}^{m}K_{i}\right)\simeq C\left(\prod_{i=1}^{n}K_{i},C\left(\prod_{i=n+1}^{m}K_{i}\right)\right)\simeq C\left(\prod_{i=1}^{n}K_{i},C(M)\right).$
Using Theorem 1.3 we can also see that
$C\left(\prod_{i=1}^{n}K_{i}\right)\simeq C\left(\prod_{i=1}^{n-1}K_{i},C(K_{n})\right)\simeq C\left(\prod_{i=1}^{n-1}K_{i},C(I_{\mathbf{A}})\right)\simeq C\left(I_{\mathbf{A}}\times\prod_{i=1}^{n-1}K_{i}\right),$
so by continuing this process further we obtain $C(\prod_{i=1}^{n}K_{i})\simeq C((I_{\mathbf{A}})^{n})$. It follows that
$C\left(\prod_{i=1}^{m}K_{i}\right)\simeq C\big((I_{\mathbf{A}})^{n}\times M\big).$
By the results of Bessaga, Pełczyński [5] and Miljutin [26] we know that $C(M)$ is isomorphic to $C(M^{\prime})$, where
$M^{\prime}\in\{\{0,1,\dotsc,n-1\}:n\in\omega\}\cup\{[0,1]\}\cup\{\omega^{\omega^{\alpha}}+1:0\leq\alpha<\omega_{1}\}.$
By Facts 8.2 and 8.3 we have
$C\big((I_{\mathbf{A}})^{n}\times M\big)\simeq C\big((I_{\mathbf{A}})^{n},C(M)\big)\simeq C\big((I_{\mathbf{A}})^{n},C(M^{\prime})\big)\simeq C\big((I_{\mathbf{A}})^{n}\times M^{\prime}\big).$
Additionally, if $C(M^{\prime})$ is finite-dimensional (i.e., $M^{\prime}$ is finite), then we have
$C\big((I_{\mathbf{A}})^{n}\times M^{\prime}\big)$ $\simeq$ $C\big((I_{\mathbf{A}})^{n-1}\times I_{\mathbf{A}}\times M^{\prime}\big)\simeq C\big((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}}\times M^{\prime})\big)$
$\simeq$ $C\big((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}})\big)\simeq C((I_{\mathbf{A}})^{n})$
due to Theorem 1.3 and the fact that the space $I_{\mathbf{A}}\times M^{\prime}$ is homeomorphic to a separable compact line.
If $C(M^{\prime})$ is isomorphic to $c_{0}$ (i.e., $M^{\prime}=\omega+1$), then
$C((I_{\mathbf{A}})^{n}\times M^{\prime})$ $\simeq$ $C((I_{\mathbf{A}})^{n-1}\times I_{\mathbf{A}}\times M^{\prime})\simeq C((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}}\times M^{\prime}))$
$\simeq$ $C((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}},C(M^{\prime})))\simeq C((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}},c_{0}))$
$\simeq$ $C((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}}))\simeq C((I_{\mathbf{A}})^{n})$
due to Lemma 8.4. ∎
9. Final remarks
9.1. Compact spaces of weight $\kappa$
Theorem 1.1 should be compared with the corollary from the Miljutin, Bessaga and Pełczyński’s classification of separable Banach spaces of the form $C(K)$ which says that there are only $\omega_{1}$ mutually nonisomorphic such spaces regardless of the relation between $\omega_{1}$ and $2^{\omega}$ (although the spaces $\ell_{p}$ for $1<p<\infty$ form a family of cardinality $2^{\omega}$ consisting of mutually nonisomorphic separable Banach spaces). Our method of proving Proposition 1.1 works only for regular cardinals. In particular, we do not know the answer to the following:
Question 1.
Suppose that $\kappa$ is a singular cardinal. Is it provable without any additional hypothesis that there are $2^{\kappa}$ mutually nonisomorphic Banach spaces of the form $C(K)$ (of any form) of density $\kappa$?
Note that the answer to the above questions is positive if we assume the Generalized Continuum Hypothesis (GCH). Indeed, Kisljakov’s classification of spaces of the form $C([0,\alpha])$ for $\alpha<\kappa^{+}$ ([19]) shows that without any additional hypothesis we have $\kappa^{+}$ mutually nonisomorphic Banach spaces of the form $C(K)$ for any uncountable cardinal $\kappa$. As GCH means that $2^{\kappa}=\kappa^{+}$ for every infinite cardinal $\kappa$, we obtain the positive answer to the above questions under GCH.
Also note that we cannot have more than $2^{\kappa}$ mutually nonisometric (and so nonisomorphic) Banach spaces of the form $C(K)$. This follows from the fact that all compact spaces of weight not exceeding $\kappa$ embed homeomorphically into $0,1]^{\kappa}$ and this space has at most $2^{\kappa}$ closed subsets as they all are complements of some unions of elements of the basis. This also implies that there are at most $2^{\kappa}$ mutually nonisometric general Banach spaces of density $\kappa$. Indeed, each such space isometrically embeds into some space of the form $C(K)$ of density $\kappa$ and there are only $2^{\kappa}$ closed subspaces of such a $C(K)$ as they are again complements of unions of some open balls around the points in the dense set of cardinality $\kappa$.
It should also be noted that for every uncountable cardinal $\kappa$ (including singular cardinals), there are $2^{\kappa}$ mutually nonhomeomorphic compact spaces of weight $\kappa$. This follows from the Stone duality and the existence of many examples of families of cardinalities $2^{\kappa}$ of mutually nonisomorphic Boolean algebras of cardinality $\kappa$ presented in [27]. In particular, it is shown in Theorem 1.3 of [27] that for every uncountable $\kappa$ there are $2^{\kappa}$ mutually nonisomorphic interval Boolean algebras. This together with Theorem 15.7 from [20] implies that there are $2^{\kappa}$ mutually nonhomeomorphic linearly ordered compact spaces for each uncountable cardinal $\kappa$.
Let us also mention that the possibility of using families of pairwise disjoint stationary subsets of $\omega_{1}$ to build $2^{\omega_{1}}$ different structures of “sizes” $\omega_{1}$ as in the proof of Theorem 1.1 has been well known. For example, this argument shows that there are many nonhomeomorphic 2-dimensional nonmetrizable manifolds (Examples 6.2 of [30]) or is used in [35] in the context of continua.
9.2. Separable compact lines
The whole proof of Theorem 1.3 relies on Baumgartner’s Theorem (Theorem 7.2). In fact, we do not need its full form, but only a weaker version stating that for two $\omega_{1}$-dense subsets $A,B$ of $(0,1)$ spaces $C(I_{A})$ and $C(I_{B})$ are isomorphic. This statement may perhaps be consistently proven under weaker assumptions, for example, that every two $\omega_{1}$-dense subsets $A,B$ of $(0,1)$ there is an order embedding of $A$ to $B$. Many interesting facts about this and similar axioms can be found in [1, Section 6].
[…]
For every separable compact line $L$ of weight $\kappa$ and every $\kappa$-dense set $B\subseteq(0,1)$ the space $C(L)$ can be embedded in $C(I_{B})$ as a complemented subspace.
Proof.
By repeating the proof of Lemma 7.6, the space $C(L)$ is isomorphic to $C(L^{\prime})\oplus C(N)$ for a 0-dimensional perfect separable compact line $L^{\prime}$ of weight $\kappa$ and some metrizable compact space $N$. But by Lemma 5.4, we have
$C(L^{\prime})\simeq C(L^{\prime})\oplus C(N)\simeq C(L).$
By Proposition 4.7, the space $L^{\prime}$ is homeomorphic to $C_{A}$ for some set $A\subseteq C$ of cardinality $\kappa$. Note that for $\kappa>\omega_{1}$ the analogue of Corollary 7.4 also holds. The set $(B\setminus C)\cup A$ is still $\kappa$-dense in $(0,1)$, so the spaces $I_{B}$ and $I_{(B\setminus C)\cup A}$ are homeomorphic. It is standard to check that $C_{A}$ is a closed subset of $I_{(B\setminus C)\cup A}$, so by Lemma 5.3 we have
$C(C_{A})|C(I_{(B\setminus C)\cup A}),$
which ends the proof, as $C(L)\simeq C(L^{\prime})\simeq C(C_{A})$ and $C(I_{(B\setminus C)\cup A})\simeq C(I_{B})$. ∎
[…]
If $\kappa$ has uncountable cofinality then, for every separable compact line $L$ of weight $\kappa$ and every $\kappa$-dense set $B\subseteq(0,1)$, the space $C(I_{B})$ can be embedded in $C(L)$ as a complemented subspace.
Proof.
By Lemma 5.3 it is enough to show that $L$ contains a topological copy of $I_{B}$. To achieve this we need again to use some ideas from the proof of Lemma 7.6.
As usual, we can assume that $L=I_{A}$, where $A$ is a subset of $(0,1)$ of size $\kappa$. Let
$U=\bigcup\{(a,b)\subseteq(0,1):a,b\in\mathbb{Q}\mbox{ and }|(a,b)\cap A|<\kappa\}.$
Clearly, $|U\cap A|<\kappa$ (by $cf(\kappa)>\omega$) and $A\setminus U$ is $\kappa$-dense in a compact set $M=I\setminus U$. We consider a separable compact line $K=M_{A\setminus U}\subseteq L$. Let $\alpha=\min\{\beta:K^{(\beta+1)}=K^{(\beta)}\}$. In the same way as in the proof of Lemma 5.3 we can check that the subspace $K^{(\alpha)}$ of $L$ is homeomorphic to $I_{B}$. ∎